OpenStudy (anonymous):

Solve by substitution : y – 3x = 5 y + x = 3

OpenStudy (anonymous):

y – 3x = 5....(i) y + x = 3......(ii) from (i) y=5+3x substitute this in (ii) 5+3x+x = 3 5+4x = 3 4x = 3-5 4x =-2 x =? y=?

OpenStudy (precal):

once you find x, sub that back into (i) or (ii) to find y